# A Rollout-Based Algorithm and Reward Function for Efficient Resource Allocation in Business Processes

### v1

@@ 4.3 Reward function
Figure 4: Example of the proposed reward function. The reward for each action is returned after the state transition happens. The reward is equal to the time passed multiplied by the cases in the process at that time.
@@ 4.4 Uniformization of the CTMDP and optimal policy
Computing the optimal policy of a CTMDP using value iteration is not straightforward due to the stochastic transition times on which the rewards depend. Therefore, we first introduce a transformation to uniformize the CTMDP into a discrete-time MDP [22]. We can then apply the value iteration algorithm to the MDP to compute the optimal policy. This policy is also optimal for the CTMDP in terms of average reward [22]. First, we choose $\tau$ to represent the time between states, such that $t\in T=\{n\tau\mid n\in\mathbb{N}_{0}\}$:
@@ 5 Evaluation
In this section, we provide the results of experiments to evaluate our algorithm and reward function. In Section 5.1, we introduce our evaluation protocol and hyperparameter tuning procedure. In Section 5.2, we demonstrate the capability of our method to learn optimal policies. Last, Section 5.3 shows the performance evaluation of existing reward functions in the literature using a PPO algorithm.
@@ 5.1 Evaluation setup
We use the business processes, or scenarios, from Middelhuis et al. [12] representing typical process flows and structures seen in business processes. The scenarios are shown in Figure 5.
@@
Figure 5: Six scenarios with two activities and resources. In each scenario, cases arrive with a rate of $\lambda=1/2$. Activity $act$ is executed by resource $r$ with rate $1/\mathbb{E}[X_{r,act}]$.
Similar to Middelhuis et al. [12], we simulate 2500 cases per episode, which terminates when all cases are completed. When evaluating our policies, both during training and evaluation, we use 300 episodes. In our results, we report the mean cycle time and 95% confidence interval (CI).
Before training and evaluating our method, we tune two hyperparameters of our algorithm and one parameter related to the CTMDP-to-MDP transformation. We tune the hyperparameters on the MDP, and the selected values are applied when training all policies, including those on the CTMDP. We tune the number of rollouts per action, $M\in{50,100,200}$, and rollout length, $N\in{25,50,100}$, and the transformation parameter $\kappa$, which determines the MDP step size $\tau=\kappa\min_{s,a}\tau_{s}(a)$. To ensure consistency between MDP and CTMDP settings, we scale the rollout length $N$ with $\tau$. For the neural network architecture, we use two layers of 128 neurons.
Through grid search, we identify the optimal hyperparameters. A value of $M=100$ provides a reliable estimate of the return, with no significant improvement seen by increasing this value. A longer rollout length of $N=100$ gives the agent a better indication of the long-term effects of rollouts while still being able to distinguish between different actions. Finally, we found that $\kappa=0.5$ yields the best results, aligning with the recommendation of Tijms [22].
@@ 5.2 Evaluation results of the proposed method
In this section, we describe the training and evaluation procedure of our proposed method in six scenarios. We train on both the MDP and CTMDP and evaluate each method on the CTMDP, as it represents the original business process. For each model, we generate 5000 random states in each iteration and perform 10 training iterations, for which we save the best model. Table 1 shows the evaluation of our method on both CTMDP and MDP compared to the optimal policy and three heuristics: Shortest Processing Time (SPT), First-In-First-Out (FIFO), and a random policy. We test for significance using a t-test and present results that are not significantly different from the optimal policy in bold.
@@
Scenario Optimal Rollout-based algorithm SPT FIFO Random
policy CTMDP MDP
Low Utilization 5.7 (0.05) 0.53% 1.24% 4.07% 5.13% 17.88%
High Utilization 18.0 (0.55) 3.78% 0.89% 9.12% 49.22% 100.56%
Slow Server 11.7 (0.23) 2.56% 1.29% 151.07% 58.58% 94.88%
Slow Downstream 10.1 (0.20) -0.30% 1.00% 46.76% -0.50% 14.93%
N-system 5.8 (0.06) 0.17% 1.55% 23.79% 1.38% 11.72%
Parallel 9.7 (0.19) 1.13% 1.55% 44.02% -2.68% 16.39%
@@ 5.3 Evaluation results of PPO with different reward functions
Each existing reward function gives a reward when case $c\in\mathcal{C}$ is completed: (a) +1, (b) $-CT_{c}$, and (c) $\frac{1}{1+CT_{c}}$. In state transition following actions that did not result in a case completion, a reward of zero is returned. A PPO agent combined with reward function (c) is considered the current state-of-the-art method for resource allocation in business processes [11, 12].
@@
Optimal Reward function Random
policy (a): +1 (b): $-CT_{c}$ (c): $\frac{1}{1+CT_{c}}$ (d): $-|C|(t-t^{\prime})$
@@
In conclusion, existing reward functions in combination with a PPO agent are not sufficient to learn competitive policies. On the one hand, the reward functions do not provide a reward signal that is good enough to guide the agent in minimizing the mean cycle time. On the other hand, the PPO algorithm cannot handle the dynamic environment in which it takes actions, leading to suboptimal policies for most business processes. Our method learns the optimal policy in all scenarios, outperforming the state-of-the-art in four scenarios and learning a similar policy in the other two scenarios.
The code and trained models used to produce our results can be found in our repository: https://anonymous.4open.science/r/BPO_rollouts-D884.
@@ 6 Conclusion
The results show that our proposed method consistently learns the optimal policy for the six evaluated business processes. In contrast, the current state-of-the-art algorithm learns the optimal policy in only two. The benefit of using rollouts is that we can directly evaluate the outcome of different actions and do not need to approximate the value function, which is typical for other DRL methods [17, 13]. Furthermore, a comparative analysis highlighted that while existing reward functions guide agents toward cycle time minimization, they can often be exploited, leading to suboptimal or unintended behaviors. In contrast, maximizing our proposed reward function ensures that the objective is minimized, eliminating the need for problem-specific reward engineering.
In this paper, we presented a method to learn optimal policies for business processes. While we can compute the optimal policy on small business processes, we cannot do so for larger processes due to state explosion. In future work, the scalability of our method to more complex processes should be explored. Furthermore, the business process with parallelism illustrated that the current state and action representation can be enhanced with case-specific features to learn better policies, which will be investigated in future research.
@@ 6.0.1 Acknowledgements
The research leading up to this paper is supported by the Dutch foundation for scientific research (NWO) under the CERTIF-AI project (grant nr. 17998).

### v2

@@ 4.3 Reward function
Figure 4: Example of the proposed reward function for three cases. At each decision step, the reward function returns a reward equal to the contribution to the total cycle time during the state transition. Waiting cases (hatched area) also incur cycle time. The reward is returned after the state transition.
@@ 4.4 Uniformization of the CTMDP and optimal policy
Computing the optimal policy of a CTMDP using value iteration is not straightforward due to the stochastic transition times on which the rewards depend. Therefore, we first introduce a transformation to uniformize the CTMDP into a discrete-time MDP [24]. We can then apply the value iteration algorithm to the MDP to compute the optimal policy. This policy is also optimal for the CTMDP in terms of average reward [24]. Note that the MDP is only required to compute the optimal policy and is not required for our algorithm presented in Section 4.2. First, we choose $\tau$ to represent the time between states, such that $t\in T=\{n\tau\mid n\in\mathbb{N}_{0}\}$:
@@ 5 Evaluation
This Section evaluates our approach. Section 5.1 introduces our evaluation protocol. Section 5.2 demonstrates the ability of our method to learn optimal policies. Section 5.3 shows the performance of different reward functions and Section 5.4 shows an evaluation on more realistically sized models.
@@ 5.1 Evaluation setup
For the evaluation, we use the scenarios from Middelhuis et al. [12] shown in Figure 5, which represent typical process patterns.
@@
Figure 5: Six scenarios with two activities and resources. In each scenario, cases arrive with a rate of $\lambda=1/2$. Activity actact is executed by resource $r$ with rate $1/\mathbb{E}[X_{r,act}]$.
Similar to Middelhuis et al. [12], we simulate 2500 cases per episode, which terminates when all cases are completed. For evaluation we use 300 episodes. Before training and evaluating our method, we tune hyperparameters on the MDP through grid search. The selected parameter values are applied when training all policies. We tune the number of rollouts per action, $M\in\{50,100,200\}$, and rollout length, $N\in\{25,50,100\}$, and the transformation parameter $\kappa\in\{0.1,0.25,0.5,1\}$, which determines the MDP step size $\tau=\kappa\min_{s,a}\tau_{s}(a)$. To ensure consistency between MDP and CTMDP settings, we scale the rollout length $N$ with $\tau$. Hyperparameter tuning leads to a value of $M=100$. A longer rollout length of $N=100$, and a $\kappa=0.5$, aligning with the recommendation of Tijms [24]. For the neural network architecture, we use two layers of 128 neurons. For each model, we generate 5000 random states in each iteration and perform 10 training iterations, of which we save the best model.
@@ 5.2 Evaluation results of the proposed method
We train on both the CTMDP and MDP versions of the six scenarios and evaluate each method on the CTMDP, as it represents the original business process. While the transformation to MDP is not required to use our proposed method, we train an agent on the MDP to investigate if the environment with more predictable rewards impacts the learning process.
Table 1 shows the results, comparing our method, trained on both CTMDP and MDP, to the optimal policy and three heuristics: Shortest Processing Time (SPT), First-In-First-Out (FIFO), and a random policy. We use a t-test ($\alpha$ = 0.05) to assess statistical significance. Results not significantly different from the optimal policy shown in bold. For each method, we show the optimality gap: the percentage difference to the optimal policy.
@@
Scenario
Optimal
policy
Our method SPT FIFO Random
CTMDP MDP
Low Utilization 5.7 (0.05) -0.41% 1.03% 4.05% 5.05% 17.78%
High Utilization 18.0 (0.55) 1.31% 2.40% 9.10% 49.23% 100.52%
Slow Server 11.7 (0.23) -0.47% 1.60% 151.14% 58.57% 94.90%
Slow Downstream 10.1 (0.20) -1.29% -0.34% 46.79% -0.51% 14.94%
N-system 5.8 (0.06) -1.33% -1.15% 23.88% 1.31% 11.71%
Parallel 9.7 (0.19) 1.56% 1.06% 44.00% -2.68% 16.41%
@@ 5.3 Evaluation results of PPO with different reward functions
Each existing reward function gives a reward when case $c\in\mathcal{C}$ is completed: (a) +1, (b) $CT_{c}$, and (c) $\frac{1}{1+CT_{c}}$. In state transition following actions that did not result in a case completion, a reward of zero is returned. A PPO agent combined with reward function (c) is considered the current state-of-the-art method for resource allocation in business processes [11, 12].
@@
The code and trained models used to produce our results can be found in our repository: https://github.com/jeroenmiddelhuis/BPO_rollouts.
@@
Optimal policy Reward function Random
(a): +1 (b): $CT_{c}$ (c): $\frac{1}{1+CT_{c}}$ (d): $|C|(t-t^{\prime})$
@@
In conclusion, existing reward functions in combination with a PPO agent are not sufficient to learn competitive policies, while our method learns the optimal policy in all scenarios, outperforming the state-of-the-art in four scenarios and learning a similar policy in the other two scenarios.
5.4 Evaluation results on composite process models
This section demonstrates our algorithm’s capability to learn policies for increasingly complex process models, even as computational complexity grows.
The computational complexity of the algorithm can be measured in terms of the speed at which it learns a good policy, which depends on: the size of the neural network, consequently translating to size of state and action space, and the number of steps between taking an action and seeing the effect of that action (i.e. the moment at which a case completes). These factors are both determined by the number of tasks and the number of resources in a process model. Therefore, we gradually increase the size of a process model and observe the effect on the learning speed.
Specifically, starting with scenario 1, we add scenarios sequentially such that the departures of one scenario are the arrivals of the next, until we have a model that has the same complexity as the composite model from [12]. We train the algorithm with the same hyperparameters only on the CTMDP, as it represents the original business process. We do not train an agent on the MDP as the results show no significant differences with different environment representations (Table 1). We removed the ‘postpone’ action, because preliminary experiments revealed that it only added noise to the training dataset, and it was rarely selected as the best action. Finally, we report on the learning speed by measuring the number of learning iterations at which our model learns a policy competitive with the heuristics.
Table 3: Mean cycle time with 95% confidence intervals for composite process models comparing our method against heuristic benchmarks, and the number of iterations required for our method to achieve performance equal to or better than the best heuristic.
Scenarios Our method Iterations SPT FIFO Random
1 5.6 (0.05) 1 5.9 (0.05) 5.9 (0.06) 6.7 (0.08)
1-2 24.1 (0.67) 2 25.0 (0.53) 32.7 (1.54) 40.9 (1.73)
1-2-3 36.7 (0.79) 1 52.7 (1.52) 47.0 (1.70) 60.9 (2.29)
1-2-3-4 50.7 (1.00) 2 71.1 (1.67) 55.8 (1.81) 72.4 (2.46)
1-2-3-4-5 57.6 (1.38) 3 81.7 (1.89) 61.4 (1.72) 76.9 (2.19)
1-2-3-4-5-6 69.6 (1.19) 6 103.1 (2.56) 70.8 (2.10) 87.6 (2.51)
Table 3 shows the results of the comparison. The table shows that for all levels of complexity, our method can learn a policy that performs at least as well as the best heuristic. While the computational complexity seems to increase exponentially, the number of learning iterations for these processes is still small and computationally manageable. In future work, further algorithmic improvements, such as those proposed by Temizöz et al. [23], and code optimization can still be applied to speed up the learning process.
@@ 6 Conclusion
The results show that our proposed method consistently learns the optimal policy for the six evaluated business processes. In contrast, the current state-of-the-art algorithm learns the optimal policy in only two. In four of five composite processes, our method learns a policy that outperforms the best heuristic, and matches it in the fifth. The benefit of using rollouts is that we can directly evaluate the outcome of different actions and do not need to approximate the value function, which is typical for other DRL methods [13, 19]. Furthermore, a comparative analysis highlighted that while existing reward functions guide agents toward cycle time minimization, they can often be exploited, leading to suboptimal or unintended behaviors. On the other hand, maximizing our proposed reward function ensures that the objective is minimized, eliminating the need for problem-specific reward engineering.
In this paper, we presented a method to learn optimal policies for business processes. While we can compute the optimal policy on small business processes, we cannot do so for larger processes due to state explosion. In future work, the ability of our method to optimize real-world processes should be explored. Furthermore, the business process with parallelism illustrated that the current state and action representation can be enhanced with case-specific features to learn better policies, which will also be investigated in future research.
@@ 6.0.1 Acknowledgements
The research leading to this paper is supported by the Dutch foundation for scientific research (NWO) under the CERTIF-AI project (nr. 17998).